The earth exerts a gravitational force of 1,9 x 1020 N on the moon. The mass of the moon is 7,4 x 1022 kg and its radius is 1,74 x 106 m.

An astronomer that weighs 20 kg on earth, lands on the moon. What would his mass be on the moon?


1
Expert's answer
2021-11-26T10:29:13-0500

"m=20 \; kg \\

g = 9.8 \; m/s^2"

weight of object on Earth "= mg = 20 \times 9.8 = 196 \;N"

Weight of object on Moon "= \frac{1}{6} \times" (weight of object on Earth)

"= \frac{1}{6} \times 196 \\

= 32.66 \;N"